# Assisting Mathematical Formalization with A Learning-based Premise Retriever

### v2

@@ Abstract
Premise selection is a crucial yet challenging step in mathematical formalization, especially for users with limited experience. Due to the lack of available formalization projects, existing approaches that leverage language models often suffer from data scarcity. In this work, we introduce an innovative method for training a premise retriever to support the formalization of mathematics. Our approach employs a BERT model to embed proof states and premises into a shared latent space. The retrieval model is trained within a contrastive learning framework and incorporates a domain-specific tokenizer along with a fine-grained similarity computation method. Experimental results show that our model is highly competitive compared to existing baselines, achieving strong performance while requiring fewer computational resources. Performance is further enhanced through the integration of a re-ranking module. To streamline the formalization process, we release a search engine (https://premise-search.com) that enables users to query Mathlib theorems directly using proof states, significantly improving accessibility and efficiency. Codes are available at https://github.com/ruc-ai4math/Premise-Retrieval.
@@ 1 Introduction
Formalized mathematics has recently attracted significant attention. It helps verify existing mathematical results, identify errors in the literature, and has the potential to accelerate the peer-review process for mathematical papers. This process involves proving natural language theorems within a strict formal logical framework. Interactive theorem provers (ITPs), or proof assistants, are commonly used for this purpose. In recent years, mathematicians have actively participated in developing large formalized theorem libraries using ITPs such as Lean [7], Coq [2], and Isabelle [25].
@@
Figure 1: (a) A schematic diagram illustrating user interactions with an ITP (e.g., Lean) given our premise retrieval model. (b) Examples of the theorem-proving process in Lean.
@@
However, the process of mathematics formalization often demands extensive experience with formalization and a high level of mathematical expertise. Therefore, there has been a sustained demand for automated auxiliary tools in the formalization field, such as premise retrievers. As illustrated in Figure 1(a), premise retrieval aims to retrieve useful premises based on the current proof state provided by the formalization platform. An effective formal theorem premise retrieval system can help mathematicians quickly identify relevant theorems, thereby accelerating the proof process. For example, LeanDojo [41] shows that a retrieval-augmented generator achieves a higher proof pass rate compared to a standalone generator without premise selection.
@@
The early premise retrieval methods leverage internal APIs of ITPs to identify applicable theorems through pattern-matching algorithms. These methods, however, rely on strict matching criteria and may miss relevant theorems. Recently, learning-based methods [41, 24] have been proposed to encode formalized theorems in a latent space and retrieve them by computing their embedding similarities. The learning-based methods can be further categorized based on the query language: natural language queries or formal proof state queries. Natural language-based retrieval [9] requires users to have a sufficient mathematical understanding of the current proof state, posing a significant challenge for beginners. In contrast, formal language-based retrieval is more user-friendly, but it often relies on fine-tuning pre-trained natural language models [41] because of the scarcity of formal proof data. As a result, the fundamental differences between formal and natural languages introduce a semantic gap, leading to suboptimal performance.
@@
The above challenges motivate us to design and learn a more effective premise retrieval system that directly leverages the formal proof state to retrieve premises, improving accuracy and usability for formalized mathematics. Specifically, our premise retrieval model consists of a context-free retrieval (CFR) module and a context-aware re-ranking (CAR) module, each of which takes BERT [8] as its backbone. The CFR module derives embeddings for premises and proof states (i.e., queries) and retrieves relevant premises based on their similarity to the input proof states. The CAR module is employed to reorder the retrieved premises, and accordingly, improves the top-$k$ recall of the retrieval results. Unlike approaches that merely fine-tune pre-trained language models, we first pre-train a BERT model from scratch based on existing formalized corpora, in which a new tokenizer is learned for formal language. Based on the pre-trained BERT, we further learn the CFR and CAR modules separately via contrastive learning [5, 13]. Experiments show that the retrieval results of our model surpass state-of-the-art models in performance while still maintaining computational efficiency. Furthermore, to evaluate the practicality of our premise retrieval model, we integrate it with an independently trained tactic generator, creating a retrieval-augmented automated theorem prover. We then assess this system using the test dataset and MiniF2F benchmark [45], demonstrating the effectiveness of our retrieval model in facilitating the proof process.
To expedite the formalization process and provide convenience for provers, we are releasing our model, along with a web search engine available to all users. Additionally, we have developed a toolkit to facilitate the setup of local databases and the deployment of the model for those working on individual projects. Our contributions can be summarized as follows:
@@
  • We deploy a theorem search engine [^1] https://premise-search.com featuring a real-time updating database, available free of charge to provers, and provide a toolkit to facilitate local model deployment for users.
@@ 2.1 Preliminaries in Formalization and Lean
Interactive theorem provers (ITPs) are software systems that provide a formal language for users to construct proofs verified by small, deterministic kernels. The process of encoding mathematical definitions, theorems, and proofs using this formal language is known as formalization. Modern ITPs, such as Isabelle [25], Coq [2], and Lean [7], are built on advances in type theory and have become sufficiently expressive to formalize complex, cutting-edge mathematics[3]. Among these, Lean stands out for its powerful automation tools and the extensive, community-driven mathematical library Mathlib, making it a preferred choice for mathematicians engaged in formalizing mathematical knowledge.
@@
While there are various styles of proofwriting, machine learning researchers commonly adopt the tactic-state transformation as an abstraction of the process. A tactic is a piece of program that modifies the current proof state. Experienced Lean users can develop custom tactics through meta-programming, utilizing Lean’s native automation tools. Pre-defined tactics, such as simp, linarith, and so on, have significantly enhanced the efficiency of formalization. [^2] The simp tactic uses lemmas and hypotheses to simplify the main goal target or non-dependent hypotheses. The linarith tactic attempts to find a contradiction between hypotheses that are linear (in)equalities. More information can be found at https://lean-lang.org/doc/reference/latest/Tactic-Proofs/Tactic-Reference/#tactic-ref.
@@ 2.2.1 Learning-based Retrieval Models
In the field of information retrieval, traditional methods like BM25 [12, 30] and TF-IDF [31] rely on term information and document statistics to match queries with relevant documents. These methods are limited in performance because they do not consider the semantic similarities between queries and documents. However, with the advancement of deep learning, researchers have developed learning-based retrieval models, often referred to as dense retrieval models [1, 16, 28, 27, 19, 20]. These models typically fall into two categories.
@@
The first approach [16, 28] encodes queries and documents into embeddings separately. Their similarity is then computed in the embedding space, and the top-ranked documents are retrieved as the final results. A key advantage of this method is that document embeddings can be precomputed, allowing the system to embed only the query during retrieval and improving computational efficiency significantly. However, this approach lacks direct interaction between queries and documents, which may limit its retrieval performance. The second approach [27, 26] concatenates the query and document as a single input to the encoder model, which then outputs a relevance score between them. This method can capture richer interactions between queries and documents, leading to improved performance. However, it incurs higher computational costs since the concatenation must be performed for each query-document pair. Recently, researchers have increasingly embraced the paradigm of fine-tuning pre-trained models for various retrieval tasks [21, 38, 18], leading to substantial improvements in performance.
2.2.2 Mathematics Premise Selection
Previous mathematics theorem selection differs from natural language information retrieval, they often need to retrieve mathematical formulas in the library. Mathematical formulas are highly structured and are typically represented using tree-based structures [42, 10]. Retrieval methods [6, 44, 32, 17, 47, 33, 43] for such formulas rely on structural searches, which identify similarities by matching substructures across various features. There are also methods that utilize neural networks to retrieve mathematical information [22, 33, 46, 23, 29, 14].
@@ 2.2.2 Mathematics Premise Selection
However, premise selection in Lean is different from the aforementioned scenario. Theorems in Lean are formalized as Lean code, which lacks the rich structural information inherent to mathematical formulas. As a result, Lean’s internal API [^3] https://loogle.lean-lang.org/ employs pattern matching to identify theorems applicable to the current proof state. This approach, however, enforces strict matching criteria, which often leads to failures in retrieving relevant theorems. In addition, applications like Moogle [^4] https://www.moogle.ai/ and leansearch [9] have been developed to enable searches based on natural language descriptions of the target query. However, they place a higher demand on the precision of natural language formulations for both the query and the formal theorems. In contrast, premise selection based on purely current formal proof state is a more straightforward approach. LeanDojo [41] employs a learning-based retrieval model that encodes the proof state and theorems independently, subsequently calculating their similarity to perform the retrieval task. However, its performance is less satisfactory, likely because they fine-tune the model pre-trained on natural language which fails to adapt well to the formal structure and syntax of Lean code.
2.2.3 AI-Assisted Formalization
Tools based on deep learning, particularly large language models (LLMs), have been developed to expedite the formalization process. These language models [41, 34, 39] assist by automating the formalization of statements, suggesting relevant premises, and generating proofs. Notably, LeanCopilot [34] was designed to enable Lean users to leverage pre-trained AI models in their formalization workflows. The Lean community is actively investigating further applications of AI across various formalization projects.
@@ 3 Proposed Method
In this section, we present the proposed retrieval framework for formal premise selection. As shown in Figure 2, this framework includes three stages:
  • Extract data from the formal mathematics library to construct the theorem corpus.
  • The context-free retrieval stage involves encoding the premises and proof states using a learning-based proof state encoder and calculating semantic similarities.
  • In the context-aware re-ranking stage, the retrieved premises are concatenated with the query and evaluated using a cross-encoder for classification and re-ranking.
@@ 3.1 Data Preparation
3.1.1 Task Definition
Let $\mathbb{M}=\{\bm{p}_{i}\}_{i=1}^{N}$ denote the library to be searched, where each $\bm{p}_{i}$ represents a theorem within the library. Since theorems are functions that accept hypotheses as input and yield proofs of specified propositions, $\bm{p}_{i}$ can be represented as $\{\Gamma_{\bm{p}_{i}},G_{\bm{p}_{i}}\}$, where $\Gamma_{\bm{p}_{i}}=[v_{1},\dots,v_{n}]_{\bm{p}_{i}}$ is the argument list of the theorem and $G_{i}$ is the output of the theorem. This representation is isomorphic to a proof state. Given the current proof state $\bm{s}$, the parameterized retrieval model $\Phi(\bm{s};\mathbb{M},\Theta)$ assigns a score to each premise and returns the top-$k$ results. Note that a proof state may contain several cases, denoted as $\bm{s}=\{\bm{s}^{i}\}_{i=1}^{|\bm{s}|}$, where $\bm{s}^{i}=\{\Gamma_{\bm{s}}^{i},G_{\bm{s}}^{i}\}$.
From human-written proofs, we can extract tactic steps $\mathbb{T}=\{t_{i}\}_{i=1}^{T}$. Each tactic step contains the proof state before and after executing the tactic, as well as the premises used in the tactic, denoted as $t_{i}=(\bm{s}_{i}^{before},\{\bm{p}_{ij}\}_{j=1}^{n_{i}},\bm{s}_{i}^{after})$, where $\bm{p}_{ij}\in\mathbb{M}$. We collect state-premise pairs from these tactics to form our training dataset. We consider the state before/after a tactic both relevant to the premises used in this step since the state before should satisfy the hypotheses of the premises, while the state after will contain patterns in the goal of the premises. Thus we can form our dataset as $\mathcal{D}=\{(\bm{s}_{i},\mathcal{P}_{i})\}_{i=1}^{|\mathcal{D}|}$, where $\bm{s}_{i}$ are distinct proof states and $\mathcal{P}_{i}$ is the set of relevant premises of $\bm{s}_{i}$.
3.1.2 Data Extraction
Several tools have been developed to extract data from a Lean project. To meet our requirements, we utilize the script from LeanDojo and incorporate additional features. We applied this pipeline to Mathlib [^5] We used tag v4.10.0 of mathlib4. https://github.com/leanprover-community/mathlib4/tree/v4.10.0. The extracted information covers all the premises and proofs in this project, along with metadata such as file dependencies. For each premise, we extract its argument list and output. The proofs are lists of tactics, with the state before/after the tactics and premises used in each step. Figure 3 illustrates the statistics related to context length and the number of premises in the dataset. The context length represents the length of $\Gamma$, while the premise number indicates how many premises are utilized in a specific tactic. It’s important to note that tactics that do not have any premises are excluded from these statistics.
@@ 3.2 Model Architecture
3.2.1 Context-Free Retrieval
In the context-free retrieval stage, we adopt a dense retrieval method [16]. This approach maps both the proof state and the theorem into a shared latent space, enabling retrieval through the computation of their cosine similarity. We design the model based on the BERT encoder architecture denoted as $f(\theta)$. For a given input text, we process it through the model and apply average pooling over the last hidden states to derive its latent representation, denoted as $f(\cdot;\theta)$.
@@ 3.2.1 Context-Free Retrieval
However, considering that some states may only correlate with either the arguments or the goal, we encode the premise arguments and goal separately, which can be denoted as $f(\Gamma_{\bm{p}},\theta)$ and $f(G_{\bm{p}},\theta)$. Accordingly, we design a fine-grained similarity between the given state $\bm{s}$ and premise $\bm{p}$, which can be formulated as
$\text{sim}(\bm{s};\bm{p})=\underbrace{\frac{f(\bm{s};\theta)}{\|f(\bm{s};\theta)\|}}_{\text{state embedding}}\cdot\underbrace{\frac{1}{2}\left(\frac{f(\Gamma_{\bm{p}};\theta)}{\|f(\Gamma_{\bm{p}};\theta)\|}+\frac{f(G_{\bm{p}};\theta)}{\|f(G_{\bm{p}};\theta)\|}\right)}_{\text{premise embedding}}.$ (3)
@@
3.2.2 Context-Aware Re-ranking
While facilitating the retrieval model and vector store enables fast and efficient queries, the lack of interaction between the state and premise may lead to lower accuracy. To address the limitations of this context-free approach, we adopt a context-aware method to re-ranking the results from first stage. To be more specific, we design the re-ranking model based on BERT architecture that follows the sequence-pair classification of [26]. The state and premise will be concatenated together by [SEP] as the re-ranking module’s input sequence, i.e.,
[…]
@@ 3.2.2 Context-Aware Re-ranking
where $\mathbf{h}_{[CLS]}$ is the embedding of the [CLS] token, $\mathbf{W}$ is the weight matrix, $\mathbf{b}$ is the bias term, and $\sigma$ represents the sigmoid activation function. For top-$k$ retrieval, we first use the retrieval model to find $k_{1}$ candidates. The re-ranking model serves as a filter and returns top-$k_{2}$ as the final result.
Note that, existing re-ranking methods [27, 26] improve retrieval performance by concatenating the state and each premise and passing them through their re-ranking models. This strategy has a high computational cost during inference due to the combining process of state and premise. As a trade-off, we apply a re-ranking model to refine the ranking of premises retrieved by the retrieval model.

### v3

# Learning an Effective Premise Retrieval Model for Efficient Mathematical Formalization

@@ Abstract
Formalized mathematics has recently garnered significant attention for its ability to assist mathematicians across various fields. Premise retrieval, as a common step in mathematical formalization, has been a challenge, particularly for inexperienced users. Existing retrieval methods that facilitate natural language queries require a certain level of mathematical expertise from users, while approaches based on formal languages (e.g., Lean) typically struggle with the scarcity of training data, hindering the training of effective and generalizable retrieval models. In this work, we introduce a novel method that leverages data extracted from Mathlib to train a lightweight and effective premise retrieval model. In particular, the proposed model embeds queries (i.e., proof state provided by Lean) and premises in a latent space, featuring a tokenizer specifically trained on formal corpora. The model is learned in a contrastive learning framework, in which a fine-grained similarity calculation method and a re-ranking module are applied to enhance the retrieval performance. Experimental results demonstrate that our model outperforms existing baselines, achieving higher accuracy while maintaining a lower computational load. We have released an open-source search engine based on our retrieval model at https://premise-search.com/. The source code and the trained model can be found at https://github.com/ruc-ai4math/Premise-Retrieval.
Machine Learning, ICML
@@ 1 Introduction
Formalized mathematics has recently attracted significant attention. It helps verify existing mathematical results, identify errors in the literature, and has the potential to accelerate the peer-review process for mathematical papers. This process involves proving natural language theorems within a strict formal logical framework. Interactive theorem provers (ITPs), or proof assistants, are commonly used for this purpose. In recent years, mathematicians have actively participated in developing large formalized theorem libraries using ITPs such as Lean (De Moura et al., 2015), Coq (Barras et al., 1997), and Isabelle (Nipkow et al., 2002).
@@
Figure 1: A schematic diagram illustrating user interactions with an ITP (e.g., Lean) given our premise retrieval model.
@@
However, the process of mathematics formalization often demands extensive experience with formalization and a high level of mathematical expertise. Therefore, there has been a sustained demand for automated auxiliary tools in the formalization field, such as premise retrievers. As illustrated in Figure 1, premise retrieval aims to retrieve useful premises based on the current proof state provided by the formalization platform. An effective premise retrieval system can help mathematicians quickly identify relevant theorems, thereby accelerating the proof process. For example, LeanDojo (Yang et al., 2024) shows that a retrieval-augmented generator achieves a higher proof pass rate compared to a standalone generator without premise selection.
@@
The early premise retrieval methods leverage internal APIs of ITPs to identify applicable theorems through pattern-matching algorithms. These methods, however, rely on strict matching criteria and may miss relevant theorems. Recently, learning-based methods (Yang et al., 2024; Mikuła et al., 2023) have been proposed to encode formalized theorems in a latent space and retrieve them by computing their embedding similarities. The learning-based methods can be further categorized based on the query language: natural language queries or formal proof state queries. Natural language-based retrieval (Gao et al., 2024) requires users to have a sufficient mathematical understanding of the current proof state, posing a significant challenge for beginners. In contrast, formal language-based retrieval is more user-friendly, but it often relies on fine-tuning pre-trained natural language models (Yang et al., 2024) because of the scarcity of formal proof data. As a result, the fundamental differences between formal and natural languages introduce a semantic gap, leading to suboptimal performance.
@@
The above challenges motivate us to design and learn a more effective premise retrieval system that directly leverages the formal proof state to retrieve premises, improving accuracy and usability for formalized mathematics. Specifically, our premise retrieval model consists of a context-free retrieval (CFR) module and a context-aware re-ranking (CAR) module, each of which takes BERT (Devlin et al., 2019) as its backbone. The CFR module derives embeddings for premises and proof states (i.e., queries) and retrieves relevant premises based on their similarity to the input proof states. The CAR module is employed to reorder the retrieved premises, and accordingly, improves the top-$k$ recall of the retrieval results. Unlike approaches that merely fine-tune pre-trained language models, we first pre-train a BERT model from scratch based on existing formalized corpora, in which a new tokenizer is learned for formal language. Based on the pre-trained BERT, we further learn the CFR and CAR modules separately via contrastive learning (Chen et al., 2020; He et al., 2020). Experiments show that the retrieval results of our model surpass state-of-the-art models in performance while still maintaining computational efficiency. Furthermore, to evaluate the practicality of our premise retrieval model, we integrate it with an independently trained tactic generator, creating a retrieval-augmented automated theorem prover. We then assess this system using the test dataset and MiniF2F benchmark (Zheng et al., 2022), demonstrating the effectiveness of our retrieval model in facilitating the proof process.
Our contributions can be summarized as follows:
@@
  • Based on our model, we have deployed a search engine featuring a real-time updating database, available free of charge to provers, and provide a toolkit to facilitate local model deployment for users.
@@ 2.1 Preliminaries in Formalization and Lean
While there are various styles of proofwriting, machine learning researchers commonly adopt the tactic-state transformation as an abstraction of the process. A tactic is a piece of program that modifies the current proof state. Experienced Lean users can develop custom tactics through meta-programming, utilizing Lean’s native automation tools. Pre-defined tactics, such as simp, linarith, and so on, have significantly enhanced the efficiency of formalization. [^1] The simp tactic uses lemmas and hypotheses to simplify the main goal target or non-dependent hypotheses. The linarith tactic attempts to find a contradiction between hypotheses that are linear (in)equalities.
@@
Figure 2: Examples of the theorem-proving process in Lean.
@@ 2.2 Related Work
Learning-based Retrieval Models In the field of information retrieval, traditional methods like BM25 (Harman, 1995; Robertson et al., 2009) and TF-IDF (Salton & Buckley, 1988) rely on term information and document statistics to match queries with relevant documents. With the advancement of deep learning, researchers have developed learning-based retrieval models, often referred to as dense retrieval models (Karpukhin et al., 2020; Qu et al., 2021). These models typically fall into two categories.
@@
The first approach (Karpukhin et al., 2020; Qu et al., 2021) encodes queries and documents into embeddings separately. Their similarity is then computed in the embedding space, and the top-ranked documents are retrieved as the final results. A key advantage of this method is that document embeddings can be precomputed, allowing the system to embed only the query during retrieval and improving computational efficiency significantly. However, this approach lacks direct interaction between queries and documents, which may limit its retrieval performance. The second approach (Qiao et al., 2019; Nogueira & Cho, 2019) concatenates the query and document as a single input to the encoder model, which then outputs a relevance score between them. This method can capture richer interactions between queries and documents, leading to improved performance. However, it incurs higher computational costs since the concatenation must be performed for each query-document pair. Recently, researchers have increasingly embraced the paradigm of fine-tuning pre-trained models for various retrieval tasks (Ma et al., 2024; Wang et al., 2024; Li et al., 2024), leading to substantial improvements in performance.
@@
Premise Selection in Lean Premises are theorems or hypotheses that can be used in a tactic. Lean’s internal API [^2] Loogle: https://loogle.lean-lang.org/ employs pattern matching to identify theorems applicable to the current proof state. This approach, however, enforces strict matching criteria, which often leads to failures in retrieving relevant theorems. In addition, applications like Moogle [^3] Moogle: https://www.moogle.ai/ and leansearch (Gao et al., 2024) have been developed to enable searches based on natural language descriptions of the target query. However, they place a higher demand on the precision of natural language formulations for both the query and the formal theorems. In contrast, premise selection based on purely current formal proof state is a more straightforward approach. LeanDojo (Yang et al., 2024) employs a learning-based retrieval model that encodes the proof state and theorems independently, subsequently calculating their similarity to perform the retrieval task.
@@ 3 Proposed Method
As shown in Figure 3, our work includes three stages: 1) Extract data from the formal mathematics library to construct the theorem corpus. 2) The context-free retrieval stage involves encoding the premises and proof states using a learning-based proof state encoder and calculating semantic similarities. 3) In the context-aware re-ranking stage, the retrieved premises are concatenated with the query and evaluated using a cross-encoder for classification and re-ranking.
@@ 3.1 Data Preparation
Let $\mathbb{M}=\{\bm{p}_{i}\}_{i=1}^{N}$ denote the library to be searched, where each $\bm{p}_{i}$ represents a theorem within the library. Since theorems are functions that accept hypotheses as input and yield proofs of specified propositions, $\bm{p}_{i}$ can be represented as $\{\Gamma_{\bm{p}_{i}},G_{\bm{p}_{i}}\}$, where $\Gamma_{\bm{p}_{i}}=[v_{1},\dots,v_{n}]_{\bm{p}_{i}}$ is the argument list of the theorem and $G_{\bm{p}_{i}}$ is the output of the theorem. This representation is isomorphic to a proof state. Given the current proof state $\bm{s}$, the parameterized retrieval model $\Phi(\bm{s};\mathbb{M},\Theta)$ assigns a score to each premise and returns the top-$k$ results. Note that a proof state may contain several cases, denoted as $\bm{s}=\{\bm{s}^{i}\}_{i=1}^{|\bm{s}|}$, where $\bm{s}^{i}=\{\Gamma_{\bm{s}}^{i},G_{\bm{s}}^{i}\}$.
From human-written proofs, we extract tactic steps $\mathbb{T}=\{t_{i}\}_{i=1}^{T}$. Each tactic step contains the proof state before and after executing the tactic, as well as the premises used in the tactic, denoted as $t_{i}=(\bm{s}_{i}^{\rm{before}},\{\bm{p}_{ij}\}_{j=1}^{n_{i}},\bm{s}_{i}^{\rm{after}})$, where $\bm{p}_{ij}\in\mathbb{M}$. We collect state-premise pairs from these tactics to form our training dataset. We consider the state before/after a tactic both relevant to the premises used in this step since the state before should satisfy the hypotheses of the premises, while the state after will contain patterns in the goal of the premises. Thus, we form our dataset as $\mathcal{D}=\{(\bm{s}_{i},\mathcal{P}_{i})\}_{i=1}^{|\mathcal{D}|}$, where $\bm{s}_{i}$ are distinct proof states and $\mathcal{P}_{i}$ is the set of relevant premises of $\bm{s}_{i}$.
@@ 3.2 Model Architecture
Context-Free Retrieval In the context-free retrieval stage, we adopt a dense retrieval method (Karpukhin et al., 2020). This approach maps both the proof state and the theorem into a shared latent space, enabling retrieval through the computation of their cosine similarity. We design the model based on the BERT encoder architecture. For a given input text, we process it through the model and apply average pooling over the last hidden states to derive its latent representation, denoted as $f(\cdot;\theta)$.
@@
However, considering that some states may only correlate with either the arguments or the goal, we encode the premise arguments and goal separately, which can be denoted as $f(\Gamma_{\bm{p}},\theta)$ and $f(G_{\bm{p}},\theta)$. Accordingly, we design a fine-grained similarity between the given state $\bm{s}$ and premise $\bm{p}$$\text{sim}(\bm{s};\bm{p})$ is formulated as
$\underbrace{\frac{f(\bm{s};\theta)}{\|f(\bm{s};\theta)\|}}_{\text{state embedding}}\cdot\underbrace{\frac{1}{2}\left(\frac{f(\Gamma_{\bm{p}};\theta)}{\|f(\Gamma_{\bm{p}};\theta)\|}+\frac{f(G_{\bm{p}};\theta)}{\|f(G_{\bm{p}};\theta)\|}\right)}_{\text{premise embedding}}.$ (3)
@@
Context-Aware Re-ranking While facilitating the retrieval model and vector store enables fast and efficient queries, the lack of interaction between the state and premise may lead to lower accuracy. To address the limitations of this context-free approach, we adopt a context-aware method to re-ranking the results from first stage. Specifically, we design the re-ranking model based on BERT architecture that follows the sequence-pair classification of (Nogueira & Cho, 2019). The state and premise will be concatenated together by [SEP] as the re-ranking module’s input sequence.
@@
where $\mathbf{h}_{[CLS]}$ is the embedding of the [CLS] token, $\mathbf{W}$ is the weight matrix, $\mathbf{b}$ is the bias term, and $\sigma$ represents the sigmoid activation function. For top-$k$ retrieval, we first use the retrieval model to find $k_{1}$ candidates. The re-ranking model serves as a filter and returns top-$k$ as the final result.
Note that, existing re-ranking methods (Qiao et al., 2019; Nogueira & Cho, 2019) improve retrieval performance by concatenating the query and each passage and passing them through their re-ranking models. This strategy has a high computational cost during inference due to the combining process of state and premise. As a trade-off, we apply a re-ranking model to refine the ranking of premises retrieved by the retrieval model.